OpenStudy (anonymous):

I'm stuck. Evaluate the polynomial for x=2. 2x^2-4x+4 When x=2

OpenStudy (anonymous):

2(2^2)-4(2) +4 = 2(4) -8 +4 = 8 -8 +4 = 4

OpenStudy (niel):

ok nvm your answer is already en route

OpenStudy (anonymous):

thanks #)